OpenStudy (anonymous):

anyone!! (5a)(6b)(2c)=

OpenStudy (anonymous):

60abc

OpenStudy (anonymous):

What are u supposed to do? simplify iT?

OpenStudy (anonymous):

5*6*2=30

OpenStudy (anonymous):

yea simplify

OpenStudy (anonymous):

Okay 5*6=30*2=60 then add the variables to the equation 60abc